OpenStudy (anonymous):

The average of three numbers is 94. The average of four different numbers is 87. What is the weighted mean of all the numbers together?

OpenStudy (anonymous):

There are 7 numbers in total. 3 of them make up the average of 94, meaning that when you put all the numbers together, the 94 counts only 34% (from 3/7) towards the final mean. Likewise, 87 counts towards, 57% which is from 4/7. When calculating weighted averages, you do the sum of the averages, times the percentage of its weight. \[\sum x _{i} w _{i}\]Where x is the average and w is the weight in percentages. http://en.wikipedia.org/wiki/Weighted_mean#Convex_combination Here's the work: 3/7 = 43% for the average of 94 4/7 = 54% for the average of 87 94(.43) + 87(.54) = 90.01 So, the weighted mean of all the numbers is about 90. Also, next time you might want to post in the Mathematics section. http://openstudy.com/study#/groups/Mathematics

OpenStudy (anonymous):

thank you very muchooooo(:

OpenStudy (anonymous):

No problem.